OpenStudy (precal):

When the radius of a circle is multiplied by 4, the area of the new circle is 40 in^3. What was the area of the original circle?

OpenStudy (solomonzelman):

In your question is the area of the new circle 40π ? \(\LARGE\color{blue}{ \bf A=πr^{2} }\) \(\LARGE\color{blue}{ \bf A=π(x+10)^{2} }\) where x is the radius of the original circle.

OpenStudy (precal):

2.5 is what I think is the solution

OpenStudy (solomonzelman):

In your question, is the area of the new circle 40π" or just 40" ?

OpenStudy (precal):

40 inches cubed

OpenStudy (precal):

I used 4r for the radius and solved for r then I took that radius and sub back into the formula and got 2.5 or 40/16

OpenStudy (solomonzelman):

I got disconnected, sorry.

OpenStudy (precal):

that's ok, I am using an ipad and having issues as well.

OpenStudy (solomonzelman):

Damn it, again !! Anyway. \(\LARGE\color{green}{ \bf 40=π(\frac{r}{4}) ^{2} }\) --> \(\LARGE\color{green}{ \bf 40=π(\frac{r^2}{16}) }\)

OpenStudy (solomonzelman):

so the first circle is 16 times smaller then the new one, the one that is just \(\LARGE\color{red}{ \bf 40=π(r) ^{2} }\)

OpenStudy (precal):

I had to switch to my laptop. Wonder if there are issues at openstudy?

OpenStudy (precal):

Thanks for your help.

OpenStudy (solomonzelman):

You welcome !